# An Automatic Graph Construction Framework based on Large Language Models for Recommendation

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
However, most of the existing GNN-based recommendation methods primarily focus on optimizing the model structures (Shi et al., 2018; Wang et al., 2023b) and learning strategies (Jiang et al., 2023; Wu et al., 2021) based on the pre-defined graphs, generally neglecting the importance of the graph construction stage. The quality of the graph structure is the foundation for the entire graph learning process, and can directly influence the model’s ability to capture the underlying relationships and patterns within the data (Qiao et al., 2018; Zhong et al., 2023). Earlier works for graph construction usually employ specific rules (e.g., click as linkage, or entity extraction) or human efforts via crowdsourcing (e.g., relational annotation for knowledge graphs). They are either too simplistic to model the sophisticated semantic signals in the recommendation data, or too labor-intensive to be scalable for large-scale scenarios.
@@
Secondly, existing works generally suffer from the construction inefficiency issue due to the massive invocations of LLMs. While LLMs provide support for in-depth semantic analysis and complex topological structure mining, their intrinsic expensive inference cost poses a significant challenge to the efficiency of graph construction algorithms. Most works instruct LLMs to infer the similarity scores between nodes in a pairwise manner (Shang and Huang, 2024), and result in a time complexity of $O(N^{2})$, which is impractical for real-world scenarios where the number of users/items $N$ can easily reach million or even billion level (Lin et al., 2023). Although several works propose to conduct downsampling (Sun et al., 2023) or heuristic pre-filtering (Zhao et al., 2024) to reduce the number of calls of LLMs, they generally sacrifice the graph quality and thereby introduce noise to the constructed graph. Therefore, it is crucial to design an efficient yet effective LLM-automated graph construction method for large-scale industrial applications.
@@ 3.2.3. Graph Construction
  • Item-Item Latent Factor Edge $e_{i-q}$ connects each item node with its corresponding set of latent factor nodes $\mathcal{K}^{i}$. The edges represent the relationship that the multifaceted attributes of each item can be semantically characterized by their $\mathcal{K}^{i}$.
@@ 4.1.1. Datasets
We conduct experiments on three public datasets. We show the dataset statistics in Table 1 and give detailed data preprocessing information in Appendix B due to page limitations.
Table 1. The datasets statistics.
BookCrossing 6,853 129,018 190,825 61,631 (3) 310,719 (5)
MovieLens-1M 6,038 3,533 545,114 9,506 (5) 7,084 (3)
Amazon-Books 6,010 8,531 77,1325 6,010 (1) 10,629 (3)
@@ 4.1.3. Backbone Models
As a model-agnostic framework, AutoGraph can be incorporated with various recommendation models by providing the graph-enhanced representations. In this paper, we select four representative models as backbones to validate the effectiveness of AutoGraph for recommendation, which generally cover four different core operators for user behavior modeling (He et al., 2023). YouTubeDNN (Covington et al., 2016) uses deep neural networks (DNNs) to model the user behavior sequences via mean pooling. MIND (Li et al., 2019) extracts multi-interest user representations with capsule networks (Sabour et al., 2017). GRU4Rec (Hidasi et al., 2015) utilizes gated recurrent units (GRUs) (Chung et al., 2014) to capture the user preference from historical sequences. SASRec (Kang and McAuley, 2018) leverages self-attention mechanisms (Vaswani et al., 2017) for user interest modeling.
@@
4.1.4. Baseline Methods
Based on the backbone models, various model-agnostic methods could be applied to promote the recommendation performance. Since AutoGraph harnesses large language models for automatic graph construction and enhancement, we select the baseline methods from two perspectives: (1) LLM-augmented methods that leverage the open-world knowledge and reasoning abilities of LLMs to enhance the performance. KAR (Xi et al., 2023b) designs specific prompts to extract user/item knowledge from LLMs, which is further encoded as additional features for recommendation models. UIST (Liu et al., 2024b) adopts the LLM-based discrete semantic tokens from vector quantization as auxiliary feature IDs. (2) Graph-enhanced methods that construct and incorporate different types of graphs for recommendation enhancement. LightGCN (He et al., 2020) conducts collaborative filtering based on the vanilla user-item interaction graph. CCGNN (Xv et al., 2023) uses NER techniques to extract phrases from item texts and incorporates them as explicit nodes and linkages for graph construction. TopoI2I (Sun et al., 2023) takes large language models as topological structure enhancers to deduce the pairwise item similarity for edge addition and removal, resulting in automatic graph construction for recommendation. We keep the same message propagator for LightGCN in the original paper (He et al., 2020), and apply GAT (Veličković et al., 2017) to CCGNN and TopoI2I for fair comparison.
Table 3. Comparison of graph construction efficiency of different methods. $N$ denotes the number of items and $E$ denotes the number of sampled candidates for each item to compute similarity with. Avg. GAUC Rel.Impr is the average GAUC relative improvement of different graph construction methods over the four backbone models on the three datasets (ML-1M, Amz-Books and BX is short for MovieLens-1M, Amazon-Books and BookCrossing respectively).
@@ 4.2. Overall Performance
  • AutoGraph generally outperforms existing graph-enhanced baseline methods by a significant margin. These methods either employ specific rules (i.e., click as linkage for LightGCN, and entity extraction for CCGNN) which fail to model the complex semantic signals, or focus on local-view pairwise comparison (i.e., TopoI2I) which is unable to capture the high-quality topological structure provided by the global view. In comparison, based on the latent factors from LLM knowledge, AutoGraph not only captures in-depth semantics in multiple facets, but also equips the graph with a global view, leading to superior performance.
@@ 4.3. Efficiency Comparison
The results are reported in Table 3. The details of evaluation settings are included in Appendix E due to page limitation. Next we provide a comprehensive evaluation of different graph construction methods:
@@
4.4. In-depth Analysis
In this section, we investigate the impact of different configurations and hyperparameters in AutoGraph  including the contribution of different metapaths to the graph, the strategy for generating semantic factors, and the number of codebook levels. The experiments are mainly conducted on the MovieLens-1M and Amazon-Books datasets with YT-DNN (abbr. of YouTubeDNN) and MIND as backbone models. Due to the page limitation, we further provide the visualization and case study of learned latent factors in Appendix F.
4.4.1. Impact of different metapaths
@@
4.4.2. The strategy for generating semantic factors
4.4.3. The number of codebook levels for graph construction
Backbone Model Strategy MovieLens-1M Amazon-Books
NG@10 HR@10 MRR GAUC NG@10 HR@10 MRR GAUC
YT-DNN LSH 0.0560 0.1045 0.0544 0.9040 0.0486 0.0956 0.0451 0.8195
HC 0.0578 0.1105 0.0557 0.9103 0.0552 0.1077 0.0501 0.8338
RQ (Ours) 0.0707 0.1221 0.0686 0.9128 0.0609 0.1173 0.0545 0.8362
MIND LSH 0.0469 0.0946 0.0461 0.9053 0.0592 0.1153 0.0516 0.8105
HC 0.0440 0.0861 0.0396 0.9021 0.0574 0.1112 0.0518 0.8106
RQ (Ours) 0.0643 0.1166 0.0622 0.9126 0.0737 0.1373 0.0656 0.8297
Backbone Model Levels MovieLens-1M Amazon-Books
NG@10 HR@10 MRR GAUC NG@10 HR@10 MRR GAUC
YT-DNN $N/A$ 0.0426 0.0906 0.0405 0.8963 0.0444 0.0845 0.0410 0.8417
0 0.0572 0.1039 0.0566 0.9080 0.0570 0.1125 0.0511 0.8347
0, 1 0.0707 0.1221 0.0686 0.9128 0.0609 0.1173 0.0545 0.8362
0, 1, 2 0.0678 0.1231 0.0656 0.9101 0.0561 0.1082 0.0510 0.8360
MIND $N/A$ 0.0435 0.0931 0.0410 0.8936 0.0532 0.1029 0.0487 0.8077
0 0.0584 0.1092 0.0566 0.9078 0.0737 0.1373 0.0656 0.8297
0, 1 0.0643 0.1166 0.0622 0.9126 0.0643 0.1203 0.0580 0.8019
0, 1, 2 0.0572 0.1078 0.0554 0.9094 0.0701 0.1355 0.0615 0.8283
We study the number of codebook levels used for graph construction and show the result in Table 6. Surprisingly, even though we train 3 levels of codebooks for quantization, including all of them for the final graph construction leads to suboptimal results, while using the first two levels generally yields the best performance.
4.5. Industrial Experiment
To evaluate the performance of AutoGraph, we conduct an online A/B test in Huawei’s online advertising platform for ten consecutive days, where hundreds of millions of impressions are generated these days. Specifically, 10% of users are randomly allocated to the experimental group, and another 10% to the control group. For the control group, the users are served by a highly optimized deep model. For the experimental group, the users are served by the same base model with AutoGraph. We utilize Huawei’s large language model PanGu (Zeng et al., 2021) to generate user and item knowledge, and assist the recommendation with graph-enhanced representations for the experimental group.
@@ 5.1. Graph-enhanced Recommendation
Over the past decade, GNN-based recommender systems (Ying et al., 2018; He et al., 2020; De Nadai et al., 2024) have become new state-of-the-art due to the power of graphs in capturing relationships (Grover and Leskovec, 2016; Xu et al., 2018; Hamilton et al., 2017; Perozzi et al., 2014). Most of the existing graph-enhanced recommendation methods focus on the optimization of model structures (Shi et al., 2018; Wang et al., 2023b) and improvement of learning strategies (Jiang et al., 2023; Wu et al., 2021) based on pre-defined graphs, while the importance of the graph construction stage is generally overlooked. Typically, earlier works usually construct graphs based on specific rules. Most of them construct a bipartite graph of users and items with click as linkage. (He et al., 2020; Jin et al., 2020; Chang et al., 2021). Other works (Xv et al., 2023; Wang et al., 2019c, a; Lin et al., 2015; Wang et al., 2018; Ji et al., 2021) further incorporate the semantic relationships of users and items. For example, CCGNN (Xv et al., 2023) leverages named entity recognition techniques to extract $N$-gram phrase of item queries as extra nodes, and construct a user-item-phrase tripartite graph. However, these rules (e.g., click as linkage, and entity extraction) can only explore the superficial semantic relationships and fall short of modeling the sophisticated semantic signals in recommendation data. Besides, there is also a line of works focusing on relational annotations (e.g., knowledge graphs) (Wang et al., 2019c, a; Ji et al., 2021). The annotations usually require significant human resources and specialized expertise, which is impractical in large-scale industrial scenarios. As a result, an automatic graph construction framework which is efficient and can explore deep semantics in recommendation deserves attention.
@@ 5.2. LLMs for Graph Construction
While large language models have been applied in numerous complex practical tasks and showcase significant potential (Bao et al., 2023; Jiang et al., 2024; Ahn et al., 2024; Lin et al., 2024b, a), research on LLMs for graph topology enhancement is still at an early stage (Xu et al., 2024; Zhang et al., 2023; Shang and Huang, 2024; Pan et al., 2024), and there is large blank especially for LLM-based graph topology enhancement in the context of recommender systems. Outside the field of recommender systems, the paradigm for LLM-based graph topology enhancement focuses on pairwise similarities of nodes. Specifically, LLMs are prompted to directly infer the similarity score between two nodes, then edges will be added or deleted based on the deduced similarities. Sun et al.  (Sun et al., 2023) describe the abstract information of two items to compare in the prompt, and incorporate in-context demonstrations (Brown et al., 2020) to infer the similarity. GraphEdit (Guo et al., 2024b) leverage instruction-tuning on LLMs to infer if two items belong to the same category, and then add or remove edges correspondingly. Other works mainly lie in LLMs for Knowledge Graph Completion (Chen et al., 2020; Hsu and Li, 2021; Zhang et al., 2023; Xu et al., 2024). They focus on discovering the possible relations of two entities, still belonging to the pairwise comparison paradigm.
However, in the realm of recommendation, the pairwise paradigm has disadvantages in both effectiveness and efficiency. First, in recommendation, the global view of the entire dataset (e.g., contextual information) is crucial for performance improvement (Yang et al., 2024a). However, the pairwise comparison is limited to local topology refinement of graphs and can hardly build a comprehensive global view due to the large scale of recommendation data and context window limitation. Second, the pairwise comparison of nodes incurs a time complexity of $O(N^{2})$, while the users, items and their attributes in recommendation easily reach the scale of millions (Li et al., 2019), making it impractical to be applied in real recommender systems.
In this paper, we mainly focus on utilizing LLMs for graph structure learning in recommender systems. To the best of our knowledge, we are the first to harness LLMs and vector quantization for graph construction with a global view in recommendation. A novel AutoGraph framework is proposed to enhance the graph structure with a global semantic insight, and demonstrates both effectiveness and efficiency in contrary to the existing pairwise comparison paradigm of LLMs for enhanced graph construction.
@@ A.3. Industrial Deployment
For existing users and items, the LLM-enhanced semantic vectors can be obtained and pre-stored offline. Then the quantization parameters can be trained using these vectors, and thus an enhanced graph can be constructed offline. The graph-enhanced representations can thereby be computed and pre-stored for online services. For new user/item nodes to be inserted or old user/item nodes to be updated, only one call of LLM per node is needed to get their respective semantic vectors. Through nearest neighbor search from the already-trained multi-level codebooks, latent factors of the incremental nodes can be assigned efficiently. Correspondingly, their graph-enhanced representations can be obtained and pre-stored. In practice, we do not need to frequently re-train the residual quantization model (i.e., DNN encoder, DNN decoder and codebooks) since it is expressive and generalized. We suggest that the residual quantization model can be updated in fixed intervals (e.g., one week) to re-fit the evolving user/item distribution, while during the interval the latest one is used.
@@
For MovieLens-1M and Amazon-Books datasets, we only retain the interactions with rating above 3 as positive interactions, while for BookCrossing dataset we retain the interactions with rating above 5.The records are sorted according to timestamp and we filter out users whose behavior history length is less than 5. Following previous works (Wang et al., 2019b; Qin et al., 2021; Yang et al., 2024c), each dataset is split using leave-one-out strategy, with the last item for testing, the second-to-last item for validation and the remaining interactions for training. For training, we rank the ground-truth next item against 50 randomly sampled negative items. For testing, we rank the ground-truth item against the full item sets on the MovieLens-1M and Amazon-Books dataset. On the BookCrossing dataset, we follow previous works (Wang et al., 2023a; Guo et al., 2024a; Nguyen et al., 2024) and randomly sample 1,000 negative items, as the full item scale is large. The maximum length of behavior sequence is set to 30.

### v2

@@ Abstract.
Graph Construction, Large Language Models, Recommender Systems
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737192isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Recommender systems
@@ 1. Introduction
However, most of the existing GNN-based recommendation methods primarily focus on optimizing the model structures (Shi et al., 2018; Wang et al., 2023b) and learning strategies (Jiang et al., 2023; Wu et al., 2021) based on the pre-defined graphs, generally neglecting the importance of the graph construction stage. The quality of the graph structure is the foundation for the entire graph learning process, and can directly influence the model’s ability to capture the underlying relationships and patterns within the data (Qiao et al., 2018; Zhong et al., 2023). Earlier works (Liu et al., 2023; Xv et al., 2023) for graph construction usually employ specific rules (e.g., click as linkage, or entity extraction) or human efforts via crowdsourcing (e.g., relational annotation for knowledge graphs). They are either too simplistic to model the sophisticated semantic signals in the recommendation data, or too labor-intensive to be scalable for large-scale scenarios.
@@
Secondly, existing works generally suffer from the construction inefficiency issue due to the massive invocations of LLMs. While LLMs provide support for in-depth semantic analysis and complex topological structure mining, their intrinsic expensive inference cost poses a significant challenge to the efficiency of graph construction algorithms. Most works instruct LLMs to infer the similarity scores between nodes in a pairwise manner (Shang and Huang, 2024), and result in a time complexity of $O(N^{2})$, which is impractical for real-world scenarios where the number of users/items $N$ can easily reach million or even billion level (Lin et al., 2023). Although several works propose to conduct downsampling (Sun et al., 2023; Wei et al., 2024) or heuristic pre-filtering (Zhao et al., 2024) to reduce the number of calls of LLMs, they generally sacrifice the graph quality and thereby introduce noise to the constructed graph. Therefore, it is crucial to design an efficient yet effective LLM-automated graph construction method for large-scale industrial applications.
@@ 3.2.3. Graph Construction
  • Item-Item Latent Factor Edge $e_{i-q}$ connects each item node with its corresponding set of latent factor nodes $\mathcal{K}^{i}$ learned. The edges represent the relationship that the multifaceted attributes of each item can be semantically characterized by their $\mathcal{K}^{i}$.
@@ 4.1.1. Datasets
We conduct experiments on three public datasets, i.e., MovieLens-1M, Amazon-Books and BookCrossing. Due to the page limitation, we show the dataset statistics and give detailed data preprocessing information in Appendix B.
@@ 4.1.3. Backbone Models
As a model-agnostic framework, AutoGraph can be incorporated with various recommendation models by providing the graph-enhanced representations. In this paper, we select four representative models as backbones to validate the effectiveness of AutoGraph for recommendation, i.e., YouTubeDNN (Covington et al., 2016), MIND (Li et al., 2019), GRU4Rec (Hidasi et al., 2015) and SASRec (Kang and McAuley, 2018). They generally cover four different core operators for user behavior modeling (He et al., 2023): deep neural networks (DNNs), capsule networks (Sabour et al., 2017), gated recurrent units (GRUs) (Chung et al., 2014) and self-attention mechanisms (Vaswani et al., 2017) respectively.
@@
Table 2. Comparison of graph construction efficiency of different methods. $N$ denotes the number of items and $E$ denotes the number of sampled candidates for each item to compute similarity with. Avg. GAUC Rel.Impr is the average GAUC relative improvement of different graph construction methods over the four backbone models on the three datasets .
@@
4.1.4. Baseline Methods
Based on the backbone models, various model-agnostic methods could be applied to promote recommendation performance. Since AutoGraph harnesses LLMs for automatic graph construction and enhancement, we select the baseline methods from two perspectives: (1) LLM-augmented methods that leverage the open-world knowledge and reasoning abilities of LLMs to enhance the performance. KAR (Xi et al., 2023b) designs specific prompts to extract user/item knowledge from LLMs, which is further encoded as additional features for recommendation models. UIST (Liu et al., 2024b) adopts the LLM-based discrete semantic tokens from vector quantization as auxiliary feature IDs. (2) Graph-enhanced methods that construct and incorporate different types of graphs for recommendation enhancement. LightGCN (He et al., 2020) conducts collaborative filtering based on the vanilla user-item interaction graph. CCGNN (Xv et al., 2023) uses NER techniques to extract phrases from item texts and incorporates them as explicit nodes and linkages for graph construction. TopoI2I (Sun et al., 2023) takes large language models as topological structure enhancers to deduce the pairwise item similarity for edge addition and removal, resulting in automatic graph construction for recommendation.
@@ 4.2. Overall Performance
  • AutoGraph generally outperforms existing graph-enhanced baseline methods by a significant margin. LightGCN and CCGNN employ specific rules and fail to model the complex semantic signals, while TopoI2I focuses on local-view pairwise comparison and is unable to capture the high-quality topological structure provided by the global view. In comparison, based on the latent factors from LLM knowledge, AutoGraph not only captures in-depth semantics in multiple facets, but also equips the graph with a global view, leading to superior performance.
@@ 4.3. Efficiency Comparison
The results are reported in Table 2. Due to page limitation, we explain more details of evaluation settings in Appendix E. Next we provide a comprehensive evaluation of different graph construction methods:
@@
4.4. Ablation Study
We investigate on the impact of different configurations and hyperparameters in AutoGraph. Due to the page limitation, here we only show experiments on contribution of different metapaths to the graph. More other experiments are shown in Appendix F.
@@ 4.4. Ablation Study
4.5. Industrial Deployment
To evaluate the performance of AutoGraph, we conduct an online A/B test in Huawei’s online advertising platform for ten consecutive days, where hundreds of millions of impressions are generated these days. Specifically, 10% of users are randomly allocated to the experimental group, and another 10% to the control group. For the control group, the users are served by a highly optimized deep model. For the experimental group, the users are served by the same base model with AutoGraph. We utilize Huawei’s large language model PanGu (Zeng et al., 2021) to generate user and item knowledge, and assist the recommendation with graph-enhanced representations for the experimental group. We offer more details and suggestions of industrial deployment of our AutoGraph framework in Appendix A.3.
@@ 5.1. Graph-enhanced Recommendation
Over the past decade, GNN-based recommender systems (Ying et al., 2018; He et al., 2020; De Nadai et al., 2024) have become new state-of-the-art due to the power of graphs in capturing relationships (Grover and Leskovec, 2016; Xu et al., 2018; Hamilton et al., 2017; Perozzi et al., 2014). Most of the existing graph-enhanced recommendation methods focus on the optimization of model structures (Shi et al., 2018; Wang et al., 2023b) and improvement of learning strategies (Jiang et al., 2023; Wu et al., 2021) based on pre-defined graphs, while the importance of the graph construction stage is generally overlooked. Typically, earlier works usually construct graphs based on specific rules. Most of them construct a bipartite graph of users and items with click as linkage (He et al., 2020; Jin et al., 2020; Chang et al., 2021). Other works (Xv et al., 2023; Wang et al., 2019c, a; Lin et al., 2015; Wang et al., 2018; Ji et al., 2021) further incorporate the semantic relationships of users and items. However, these designed rules (e.g., click as linkage, and entity extraction (Xv et al., 2023)) can only explore the superficial semantic relationships and fall short of modeling the sophisticated semantic signals in recommendation data. Besides, there is also a line of works focusing on relational annotations (e.g., knowledge graphs) (Wang et al., 2019c, a; Ji et al., 2021). The annotations usually require significant human resources and specialized expertise, which is impractical in large-scale industrial scenarios. As a result, it is meaningful to design an automatic graph construction framework which can explore deep semantics and is efficient for large-scale industrial applications.
@@ 5.2. LLMs for Graph Construction
While large language models have been applied in numerous complex practical tasks and showcase significant potential (Bao et al., 2023; Jiang et al., 2024; Ahn et al., 2024; Lin et al., 2024b, a; Damianou et al., 2024), research on LLMs for graph topology enhancement is still at an early stage (Xu et al., 2024; Zhang et al., 2023; Shang and Huang, 2024; Pan et al., 2024), and there is large blank especially for LLM-based graph topology enhancement in recommender systems. Outside the field of recommender systems, the paradigm for LLM-based graph topology enhancement focuses on pairwise similarities of nodes. Specifically, LLMs are prompted to directly infer the similarity score between two nodes with in-context learning strategy (Sun et al., 2023; Brown et al., 2020) or instruction tuning strategy (Guo et al., 2024b). Then edges will be added or deleted based on the deduced similarities. Other works mainly lie in LLMs for Knowledge Graph Completion (Chen et al., 2020; Hsu and Li, 2021; Zhang et al., 2023; Xu et al., 2024). They focus on discovering the possible relations of two entities, still belonging to the pairwise comparison paradigm.
However, the pairwise paradigm has disadvantages in both effectiveness and efficiency in recommendation. First, the global view of the entire dataset (e.g., contextual information) is crucial for performance improvement (Yang et al., 2024a), while the pairwise comparison is limited to local topology refinement of graphs and can hardly build a comprehensive global view, due to the large scale of recommendation data and context window limitation of LLMs. Second, the pairwise comparison of nodes incurs a time complexity of $O(N^{2})$, while the users, items and their attributes in recommendation easily reach the scale of millions (Li et al., 2019), making it impractical to be applied in real recommender systems.
In this paper, we mainly focus on how to effectively and efficiently construct graphs with LLMs on large-scale recommendation data for industrial applications. To the best of our knowledge, we are the first to harness LLMs and vector quantization for graph construction with a global view in recommendation. A novel AutoGraph framework is proposed to enhance the graph structure with a global semantic insight, and demonstrates both effectiveness and efficiency in contrary to the existing pairwise comparison paradigm of LLMs for enhanced graph construction.
@@ 6. Conclusion
Acknowledgements.
The Shanghai Jiao Tong University team is partially supported by National Key R&D Program of China (2022ZD0114804), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and National Natural Science Foundation of China (624B2096, 62322603, 62177033). The work is sponsored by Huawei Innovation Research Program. We thank MindSpore (min, 2020) for the partial support of this work, which is a new deep learning computing framework.
@@ A.3. Industrial Deployment
The new introduced nodes of AutoGraph over the vanilla user-item graphs are the latent factor nodes, whose numbers are equal to the number of quantization codebook vectors. In our practice, thousand-level codebook vectors are enough to explore meaningful semantics. The number of new nodes is much smaller than the size of users and items, which easily reaches million level. And the number of edges is controllable by neighborhood sampling, leading to graphs of appropriate sizes. Next, we provide more details of our deployment strategy.
We mainly leverage AutoGraph as a plug-and-play graph-enhanced representation generator, which is compatible with existing recommender system architectures through an offline pre-storage strategy. The core process can be divided into three stages:
  • Existing Node Processing: Offline pre-store LLM-enhanced semantic vectors, train multi-level quantization codebooks to construct enhanced graphs, and synchronize the pre-stored graph-enhanced representations for online use.
  • Incremental Node Processing: For newly added/updated nodes, a single call to the LLM is made to obtain the semantic vector. A codebook nearest neighbor search is used to quickly assign latent factors and generate pre-stored graph-enhanced representations.
  • Fast-slow Model Update: The main recommendation model is frequently refreshed, while we do not need to frequently re-train the residual quantization model (i.e., encoder, decoder and codebooks) since it is expressive and generalized. We suggest that the quantization model can be updated in fixed intervals (e.g., one week) to re-fit the evolving user/item distribution, while during the interval the latest one is used.
This solution maintains the efficiency of the online service by only adding negligible read overhead for pre-stored representations, avoiding the burden of real-time graph computation. By using offline asynchronous processing and periodic model updates, it strikes a balance between representation quality and system performance.
@@
For MovieLens-1M and Amazon-Books datasets, we only retain the interactions with rating above 3 as positive interactions, while for BookCrossing dataset we retain the interactions with rating above 5.The records are sorted according to timestamp and we filter out users whose behavior history length is less than 5. Following previous works (Wang et al., 2019b; Qin et al., 2021; Yang et al., 2024c), each dataset is split using leave-one-out strategy, with the last item for testing, the second-to-last item for validation and the remaining interactions for training. For training, we rank the ground-truth next item against 50 randomly sampled negative items. For testing, we rank the ground-truth item against the full item sets on the MovieLens-1M and Amazon-Books dataset. On the BookCrossing dataset, we follow previous works (Wang et al., 2023a; Guo et al., 2024a; Nguyen et al., 2024) and randomly sample 1,000 negative items, as the full item scale is large. The maximum length of behavior sequence is set to 30. The dataset statistics are shown in Table 4.
Table 4. The datasets statistics.
BookCrossing 6,853 129,018 190,825 61,631 (3) 310,719 (5)
MovieLens-1M 6,038 3,533 545,114 9,506 (5) 7,084 (3)
Amazon-Books 6,010 8,531 77,1325 6,010 (1) 10,629 (3)